# CircleZ : Reliable Photometric redshifts for AGN computed using only photometry from Legacy Survey Imaging for DESI

### v1

@@ Abstract
Methods. We present the results from CircleZ, a machine-learning algorithm based on a Fully Connected Neural Network. CircleZ is built on a training sample of 14,000 X-ray-detected AGN and utilizes multi-aperture photometry, mapping the light distribution of the sources.
Results. The accuracy ($\sigma_{\rm NMAD}$) and the fraction of outliers ($\eta$) reached in a test sample of 2913 AGN is 0.067 and 11.6%, respectively. The results are comparable to or better than those obtained previously for the same field but with much less effort. We further tested the stability of the results by computing the photometric redshifts for the sources detected in CSC2 and Chandra-COSMOS Legacy, reaching comparable accuracy as in eFEDS when limiting the magnitude of the counterparts with respect to the depth of LS10.
@@
giant planet formation – $\kappa$-mechanism – stability of gas spheres
@@ 1 Introduction
It is now accepted that in their lifetime, most (of the massive) galaxies will go through phase(s) during which the supermassive black hole (SMBH) that they host is accreting at a significant fraction of its Eddington limit (e.g., Madau & Dickinson, 2014). While simulations show that feedback from these Active Galactic Nuclei is necessary to regulate star-formation in their host galaxies (e.g., Di Matteo et al., 2005; Croton et al., 2006), observationally, the interplay between the host and the SMBH is, however, still a matter of study. Further progress requires that diverse samples of AGNs are studied in detail, so accurate redshift estimates are necessary. While the number of multi-slit/multifiber spectrographs mounted on large telescopes is ever increasing (e.g., SDSS:Kollmeier et al 2024, in prep; DESI:Levi et al. (2013); PFS:Takada et al. (2014); Moons:Cirasuolo et al. (2011); 4MOST: de Jong (2019)) thus raising the completeness of the spectroscopic follow-up of AGN, the urgent need to develop algorithms for the computation of reliable photometric redshifts (photo-z) for these type of objects remains; deep photometric surveys like Euclid (Mellier et al., 2024, in prep) and LSST (Ivezić et al., 2019a) will provide huge samples of AGN at a speed and depth that will not be matched by the new spectrographs!. Photo-zs for AGN are challenging because the relative host/nucleus contribution to the emission is unknown a priori and even a modest contribution to the host galaxy SEDs by the continuum emission from the AGN hides the features used for identifying the redshift. The net result is a large fraction of sources having multiple possible redshifts (degeneracies).
@@
Table 1: Modified version of the table presented in Ananna et al. (2017), which lists accuracy and the fraction of outliers obtained for X-ray surveys at different X-ray depths in the 0.5-2 keV (0.5-10 keV for XMM-SERVS), area coverage, type of photometry available (Narrow, Intermediate, Broad band photometry) and Wavelength coverage (UV, Optical, Near-Infrared, Mid-Infrared).
@@
XMM-SERVS $1\times 10^{-14}$ 4.5 B, Uv-O-NIR-MIR 0.06 18% SED fitting
@@
XMM-SERVS $1.3\times 10^{-14}$ 3.0 B, Uv-O-NIR-MIR 0.06 20% SED fitting
@@
In fact, in the computation of photo-zs for the about 22000 AGN detected by eROSITA (Predehl et al., 2021; Merloni et al., 2024) in eFEDS Salvato et al. (2022) already pointed out how the inhomogeneity of the data quality drastically affects the reliability of the results [^1] See section 7.2 of (Salvato et al., 2022), and Table 5 of this paper, where the quality of the photo-zs in the area within and outside the homogenised optical-NIR photometry from KiDS (Kuijken et al., 2019) is compared. Using the traditional method, for eRASS:8 in the best case scenario, we expect results not better than an accuracy of 0.07 and a fraction of outliers of 25%, after an extenuating effort of homogeneisation of the photometry from various surveys covering from UV to Mid-infrared. Does this showcase significant issues for computing photo-zs for AGN detected by the eROSITA all-sky surveys?
@@
We have already mentioned that the extension of the sources in optical images provides a prior on the possible redshift solutions (Salvato et al., 2009). We can go beyond the differentiation between extended and point-like sources using Machine Learning. The feature analysis of all parameters available in the SDSS DR9 (Ahn et al., 2012) and SDSS DR7 (Abazajian et al., 2009) of spectroscopically confirmed AGN and galaxies presented in (D’Isanto et al., 2018), indicated that the photo-zs quality could have been improved (e.g., smaller bias, smaller root mean square error; RMSE) for the photo-zs computed using fluxes and colors, by considering features such as Petrosian magnitudes [^2] Flux measured within a circular aperture whose radius is defined by the shape of the azimuthally averaged light profile. In particular, what was found is that ratios between magnitudes computed assuming different models (point source, exponential, de Vaoucouler, or the combination of exponential and de Vaucouler) carry relevant information. Similarly, in Nishizawa et al. (2020), the computation of photo-zs for extragalactic sources detected in the HSC survey (Aihara et al., 2018) was carried out with Machine Learning relying solely on the optical photometry measured at 1.3 and 1.5 arcsecond apertures, computed assuming a PSF, an exponential, and a de Vaucouler profile. The accuracy reached by this approach (although limited to the parameter space covered by their training sample) was comparable to or even better than the results obtained via SED fitting after combining optical data with UV, near-infrared, and mid-infrared data (see Salvato et al., 2022, for a direct comparison). Recent works are also trying to compute the photo-zs using the images directly (e.g., Schuldt et al., 2021, Roster et al., 2024, in prep). All these facts indicate value when considering the morphology, the size, and, in particular, the light distribution (e.g., in the form of an azimuth profile) when computing the photo-zs. This is because all these quantities are, at least partially, redshift-dependent. In particular, unlike the single total model flux per each photometric band (which is what most photometric surveys provide), the fluxes within different apertures and/or assuming different model profiles of light distribution (exponential, Sersic, Point-like, etc.) carry with them information on the host/nucleus relative contribution.
@@
The latest Legacy Survey imaging for DESI Data Release (LS10 Dey et al., 2019), available since January 2023, and a newly available large sample of spectroscopically confirmed AGN (see Section 3) allowed us to test this hypothesis further. Here, we present the results from CircleZ, a machine learning algorithm based on a fully-connected neural network (FCCN) that reliably computes photo-zs for AGN using only the data available from Legacy Survey without adding additional information from external surveys at different wavelengths. More than in the algorithm (very similar to DNNZ, Nishizawa et al., 2020), the innovation here is on the use of new features (colors within apertures and the quality of the fit to the azimuth profile assuming different galaxy profile models) that, just as total fluxes, themselves correlate with redshift. The increase in useful features allows us to break the degeneracy in the redshift solution typical of photo-zs for AGN computed with a limited number of photometric bands. We demonstrate this by recomputing the photo-zs of eROSITA/eFEDS and comparing the results previously obtained via SED fitting with LePHARE(Arnouts et al., 1999; Ilbert et al., 2006; Salvato et al., 2009) and via machine learning with DNNZ (Salvato et al., 2022). In addition, we tested the quality of the photo-zs for AGN selected in the Chandra Source Catalog Release 2 (CSC2666https://cxc.cfa.harvard.edu/csc/ and Legacy Chandra-COSMOS (a deep pencil-beam survey) for those sources with secure spectroscopic redshift. We show that for AGN with the counterparts of brightness typical of LS10, the photo-zs are reliable, suggesting that the method could also be adopted for fainter sources, granted that features and training samples are available. This is relevant for surveys such as Euclid (Mellier et al. 2024, in prep) and LSST (Ivezić et al., 2019b).
@@ 2 The Photometric Data
Fortunately, the Legacy imaging Survey for DESI (Dey et al., 2019) data release 10 (LS10) provides registered, background-subtracted, and calibrated ”all-sky” photometry, covering almost entirely the eROSITA-DE footprint. Despite the coverage being limited to 4 optical bands (see Fig 1) and WISE, it provides photometry that goes beyond the total flux measured assuming a galaxy profile, and the question is whether access to these features is sufficient to estimate reliable photo-zs for AGN, without the need of combining the data with Galex in UV (Bianchi, 2014) and VHS(McMahon et al., 2013) or UKIDSS(Lawrence et al., 2007) in NIR, thus simplifying the work dramatically. Interested readers can refer to the ”Content” section at https://www.legacysurvey.org/dr10/description/ for comprehensive information on the survey data, processing details, and coverage. In the following sections, we explore the features considered for CircleZ.
@@ 2.1 Going beyond total, model fluxes
LS10 is composed of two distinct parts, LS10-North and LS10-South. LS10-north is obtained by combining two optical (g,r,z) surveys (Beijing-Arizona Sky Survey–BASS [^8] https://www.legacysurvey.org/bass/; Mayall z-band Legacy Survey– MzLS [^9] https://www.legacysurvey.org/mzls/ in addition to the data obtained for DESI [^10] https://www.desi.lbl.gov/the-desi-survey/ with the Dark Energy Camera (DECam) at the CTIO Blanco 4-m telescope. It corresponds to the northern portion of the Data Release 9 (Dey et al., 2019). LS10-South, compared to the Data Release 9, has deeper data taken with DECam in g,r,i,z. For the rest of the paper, we will refer to LS10-South as LS10, as the area covered by the i band (see Figure 1), which almost completely maps the eROSITA-DE area. In particular, the data include all the observations taken within the DeROSITAs program (P.I. A. Zenteno; Zenteno et al. in prep). For each source detected in the four optical bands, LS10111111The complete list of parameters and their description are available at https://www.legacysurvey.org/dr10/catalogs/ provides PSF-forced photometry (and related errors) in the optical g,r,i,z and mid-infrared (W1, W2, W3 and W4) from all imaging through year 7 of NEOWISE-Reactivation, computed in multiple apertures (0.5, 0.75, 1.0, 1.5, 2.0, 3.5, 5.0, 7.0 arcseconds radii in the optical and 3, 5, 7, 9, 11 arcseconds radii in the mid-infrared). It also provides the total flux for each band, assuming the best model (indicated in the column TYPE: PSF”=stellar, ”REX”=”round exponential galaxy”, ”DEV”=deVauc, ”EXP”=exponential, ”SER”=Sersic profile) fitted to the azimuth profile in the r band, the quality of the fit for each model, together with an extensive catalog of observational attributes and flags indicating problematic observations.
@@
Figure 2 provides a visual representation of the features used in our analysis and demonstrates the method’s strength. The figure shows the RGB image of a normal galaxy, a local AGN, and a QSO. In the Second row, the azimuth profile of the sources is shown, with overplotted available model profiles (PSF, REX, EXP, DEV, and SER). The quality of the fits is encapsulated in the column ”dchisq” of the catalog, while the best-fitting model is used to classify the source, specified in the column ”TYPE”. In these examples, the galaxy on the left can be equally well-fitted by the ”DEV” and ”SER” profiles, with ”SER” being slightly better. For the AGN, in the middle panel, none of the models is a good profile representation, with ”EXP” being selected. This bad fitting would reflect in the quality of the photo-zs computed with algorithms that use model_flux photometry only. For each source, the flux and its residual, after removing the best model, is measured within the apertures shown in the third row. Within each aperture, all colors are computed, and the last row shows how the profiles of the colors change with the distance from the center, depending on the source type. This dependence on the size is implicitly used to compute the photo-zs correctly. For the interested reader, the importance of these features and their correlation with the redshifts are further discussed in the Appendix A.
@@
dchisq121212more details are available at https://www.legacysurvey.org/dr10/catalogs/: Difference in $\chi^{2}$ between successively more-complex model fits: PSF, REX, DEV, EXP, SER. The difference is versus no source.
@@
Figure 2: Top: Optical RGB image of an inactive galaxy (Left) and a low-redshift X-ray detected AGN (Middle) and a QSO (Right). Second row: azimuth profile of the sources, with overplotted available model profiles (PSF, REX, EXP, DEV, and SER). Third row: the apertures within which ap_flux[band] is measured. Fourth row: the residual image after subtracting the best model. Last row: within each aperture, all colors are computed; the figures show how the profiles of the colors change with the distance from the center, depending on the source type (see more details in the main text.
@@ 3.1 Construction of the sample for training
The counterparts to the training sample are then matched in Right Ascension and Declination (max separation of 1 arcsecond) with a compilation of spectroscopic redshifts available within the eROSITA collaboration. At this stage, the compilation is a mere list of catalogs from the literature collated. It encompasses approximately 13 million celestial sources, thus including duplicates and contradicting redshift measurements (see Kluge et al., submitted for more details.). Additionally, we extend the compilation by incorporating QSOs with reliable redshifts from the low-resolution spectra in Gaia DR3 (flagQSOC =0 Gaia Collaboration et al., 2023). For some sources, multiple spectroscopic matches are available, with discordant results. This can occur, for example, due to wrong line identification in low signal-to-noise spectra. Here, we prioritize redshifts from eFEDS (Merloni et al., 2024 in prep), given that most have all been visually inspected. Alternatively, a single entry is randomly selected, provided that the difference in spectroscopic redshift between the smallest and largest redshift entry $\Delta z$ does not exceed $\Delta z$ 0.05. Objects failing to meet this criterion ( $\leq 0.3\%$ of cases) are excluded from consideration to avoid the accidental assignment of incorrect redshift values to objects in the training sample. Objects with spectroscopic matches falling within the redshift range of $z\leq$ 0.002 are assumed to be galactic and, therefore, excluded from any further consideration. The compilation is enriched by including 152 proprietary entries from SDSS-V (Kollmeier et al., 2024, in prep) (SPECPRIMARY == 1, ZWARNING == 0). Within SDSS-V, the Black Hole Mapper (BHM, Anderson et al., 2024, in prep) is one of the primary surveys, specializing in long-term, time-domain investigations of AGN and the optical characterization of eROSITA X-rays sources detected during the initial 1.5 years of the all-sky survey. This part of the training sample is represented in the histogram of Figure 3 with a solid violet line.

### v2

@@ Abstract
Methods. We present the results from CircleZ, a machine-learning algorithm based on a Fully Connected Neural Network. CircleZ is built on a training sample of 14,000 X-ray-detected AGN and utilises multi-aperture photometry, mapping the light distribution of the sources.
Results. The accuracy ($\sigma_{\rm NMAD}$) and the fraction of outliers ($\eta$) reached in a test sample of 2913 AGN is 0.067 and 11.6%, respectively. The results are comparable to or better than those obtained previously for the same field but with much less effort. We further tested the stability of the results by computing the photometric redshifts for the sources detected in CSC2 and Chandra-COSMOS Legacy, reaching comparable accuracy as in eFEDS when limiting the magnitude of the counterparts to the depth of LS10.
@@
Galaxies: active – Galaxies: evolution – Surveys – Methods: statistical– Techniques: photometric
@@ 1 Introduction
It is now accepted that in their lifetime, most (of the massive) galaxies will go through phase(s) during which the supermassive black hole (SMBH) that they host is accreting at a significant fraction of its Eddington limit (e.g., Madau & Dickinson, 2014). While simulations show that feedback from these Active Galactic Nuclei is necessary to regulate star-formation in their host galaxies (e.g., Di Matteo et al., 2005; Croton et al., 2006), observationally, the interplay between the host and the SMBH is, however, still a matter of study. Further progress requires that diverse samples of AGNs are studied in detail, so accurate redshift estimates are necessary. While the number of multi-slit/multifiber spectrographs mounted on large telescopes is ever increasing (e.g., SDSS: Kollmeier et al., in prep; DESI:Levi et al. (2013); PFS:Takada et al. (2014); Moons:Cirasuolo et al. (2011); 4MOST: de Jong (2019)) thus raising the completeness of the spectroscopic follow-up of AGN, the urgent need to develop algorithms for the computation of reliable photometric redshifts (photo-z) for these type of objects remains; deep photometric surveys like Euclid (Euclid Collaboration et al., 2024) and LSST (Ivezić et al., 2019a) will provide huge samples of AGN at a speed and depth that will not be matched by the new spectrographs!. Photo-zs for AGN are challenging because the relative host/nucleus contribution to the emission is unknown a priori and even a modest contribution to the host galaxy SEDs by the continuum emission from the AGN hides the features used for identifying the redshift. The net result is a large fraction of sources having multiple possible redshifts (degeneracies).
@@
Table 1: Accuracy and fraction of outliers obtained for X-ray surveys at different X-ray depths
@@
XMM-SERVS $1\times 10^{-14}$ 4.5 B, Uv-O-NIR-MIR 0.04 9.2% SED fitting
@@
XMM-SERVS $1.3\times 10^{-14}$ 3.0 B, Uv-O-NIR-MIR 0.04 9.6% SED fitting
@@
 [^1] Modified version of the table presented in Ananna et al. (2017). The X-ray flux is measured in the 0.5-2 keV range (0.5-10 keV for XMM-SERVS). In addition, we report the area coverage, type of photometry available (Narrow, Intermediate, Broad band photometry), and wavelength coverage (UV, Optical, Near-Infrared, Mid-Infrared).
In fact, in the computation of photo-zs for the about 22000 AGN detected by eROSITA (Predehl et al., 2021; Merloni et al., 2024) in eFEDS Salvato et al. (2022) already pointed out how the inhomogeneity of the data quality drastically affects the reliability of the results [^2] See section 7.2 of (Salvato et al., 2022), and Table 23 of this paper, where the quality of the photo-zs in the area within and outside the homogenized optical-NIR photometry from KiDS (Kuijken et al., 2019) is compared. Using the traditional method, for eRASS:8 in the best case scenario, we expect results not better than an accuracy of 0.07 and a fraction of outliers of 25%, after an extenuating effort of homogenization of the photometry from various surveys covering from UV to Mid-infrared. Does this showcase significant issues for computing photo-zs for AGN detected by the eROSITA all-sky surveys?
@@
We have already mentioned that the extension of the sources in optical images provides a prior on the possible redshift solutions (Salvato et al., 2009). We can go beyond the differentiation between extended and point-like sources using Machine Learning. The feature analysis of all parameters available in the SDSS DR9 (Ahn et al., 2012) and SDSS DR7 (Abazajian et al., 2009) of spectroscopically confirmed AGN and galaxies presented in (D’Isanto et al., 2018), indicated that the photo-zs quality could have been improved (e.g., smaller bias, smaller root mean square error; RMSE) for the photo-zs computed using fluxes and colors, by considering features such as Petrosian magnitudes [^3] Flux measured within a circular aperture whose radius is defined by the shape of the azimuthally averaged light profile. In particular, what was found is that ratios between magnitudes computed assuming different models (point source, exponential, de Vaoucouler, or the combination of exponential and de Vaucouler) carry relevant information. Similarly, in Nishizawa et al. (2020), the computation of photo-zs for extragalactic sources detected in the HSC survey (Aihara et al., 2018) was carried out with Machine Learning relying solely on the optical photometry measured at 1.3 and 1.5 arcsecond apertures, computed assuming a PSF, an exponential, and a de Vaucouler profile. The accuracy reached by this approach (although limited to the parameter space covered by their training sample) was comparable to or even better than the results obtained via SED fitting after combining optical data with UV, near-infrared, and mid-infrared data (see Salvato et al., 2022, for a direct comparison). Recent works are also trying to compute the photo-zs using the images directly (e.g., Schuldt et al., 2021, Roster et al., in prep). All these facts indicate value when considering the morphology, the size, and, in particular, the light distribution (e.g., in the form of an azimuth profile) when computing the photo-zs. This is because all these quantities are, at least partially, redshift-dependent. In particular, unlike the single total model flux per each photometric band (which is what most photometric surveys provide), the fluxes within different apertures and/or assuming different model profiles of light distribution (exponential, Sersic, Point-like, etc.) carry with them information on the host/nucleus relative contribution.
@@
The latest Legacy Survey imaging for DESI Data Release (LS10 Dey et al., 2019), available since January 2023, and a newly available large sample of spectroscopically confirmed AGN (see Section 3) allowed us to test this hypothesis further. Here, we present the results from CircleZ, a machine learning algorithm based on a fully-connected neural network (FCCN) that reliably computes photo-zs for AGN using only the data available from Legacy Survey without adding additional information from external surveys at different wavelengths. More than in the algorithm (very similar to DNNZ; Nishizawa et al., 2020), the innovation here is on the use of new features (colors within apertures and the quality of the fit to the azimuth profile assuming different galaxy profile models) that, just as total fluxes, themselves correlate with redshift. The increase in useful features allows us to break the degeneracy in the redshift solution typical of photo-zs for AGN computed with a limited number of photometric bands. We demonstrate this by recomputing the photo-zs of eROSITA/eFEDS and comparing the results previously obtained via SED fitting with LePHARE(Arnouts et al., 1999; Ilbert et al., 2006; Salvato et al., 2009) and via machine learning with DNNZ (Salvato et al., 2022). In addition, we tested the quality of the photo-zs for AGN selected in the Chandra Source Catalog Release 2 (CSC2777https://cxc.cfa.harvard.edu/csc/ and Legacy Chandra-COSMOS (a deep pencil-beam survey) for those sources with secure spectroscopic redshift. We show that for AGN with the counterparts of brightness typical of LS10, the photo-zs are reliable, suggesting that the method could also be adopted for fainter sources, granted that features and training samples are available. This is relevant for surveys such as Euclid (Euclid Collaboration et al., 2024) and LSST (Ivezić et al., 2019b).
@@ 2 The Photometric Data
Fortunately, the Legacy imaging Survey for DESI (Dey et al., 2019) data release 10 (LS10) provides registered, background-subtracted, and calibrated ”all-sky” photometry, covering almost entirely the eROSITA-DE footprint. Despite the coverage being limited to 4 optical bands (see Fig 1) and WISE, it provides photometry that goes beyond the total flux measured assuming a galaxy profile, and the question is whether access to these features is sufficient to estimate reliable photo-zs for AGN, without the need of combining the data with Galex in UV (Bianchi, 2014) and VHS(McMahon et al., 2013) or UKIDSS(Lawrence et al., 2007) in NIR, thus simplifying the work dramatically. Interested readers can refer to the ”Content” section of LS10888https://www.legacysurvey.org/dr10/description/ for comprehensive information on the survey data, processing details, and coverage. In the following sections, we explore the features considered for CircleZ.
@@ 2.1 Going beyond total, model fluxes
LS10 is composed of two distinct parts, LS10-North and LS10-South. LS10-north is obtained by combining two optical (g,r,z) surveys (Beijing-Arizona Sky Survey–BASS [^10] https://www.legacysurvey.org/bass/; Mayall z-band Legacy Survey– MzLS [^11] https://www.legacysurvey.org/mzls/ in addition to the data obtained for DESI [^12] https://www.desi.lbl.gov/the-desi-survey/ with the Dark Energy Camera (DECam) at the CTIO Blanco 4-m telescope. It corresponds to the northern portion of the Data Release 9 (Dey et al., 2019). LS10-South, compared to the Data Release 9, has deeper data taken with DECam in g,r,i,z. For the rest of the paper, we will refer to LS10-South as LS10, as the area covered by the i band (see Figure 1), which almost completely maps the eROSITA-DE area. In particular, the data include all the observations taken within the DeROSITAs program (P.I. A. Zenteno; Zenteno et al., in prep). For each source detected in the four optical bands, LS10131313The complete list of parameters and their description are available at https://www.legacysurvey.org/dr10/catalogs/ provides PSF-forced photometry (and related errors) in the optical g,r,i,z and mid-infrared (W1, W2, W3 and W4) from all imaging through year 7 of NEOWISE-Reactivation, computed in multiple apertures (0.5, 0.75, 1.0, 1.5, 2.0, 3.5, 5.0, 7.0 arcsecond radii in the optical and 3, 5, 7, 9, 11 arcseconds radii in the mid-infrared). It also provides the total flux for each band, assuming the best model (indicated in the column TYPE: PSF”=stellar, ”REX”=”round exponential galaxy”, ”DEV”=deVauc, ”EXP”=exponential, ”SER”=Sersic profile) fitted to the azimuth profile in the r band, the quality of the fit for each model, together with an extensive catalog of observational attributes and flags indicating problematic observations.
@@
Figure 2 provides a visual representation of the features used in our analysis and demonstrates the method’s strength. The figure shows the RGB image of a normal galaxy, a local AGN, and a QSO. In the Second row, the azimuth profile of the sources is shown, with over-plotted available model profiles (PSF, REX, EXP, DEV, and SER). The quality of the fits is encapsulated in the column ”dchisq” of the catalog, while the best-fitting model is used to classify the source, specified in the column ”TYPE”. In these examples, the galaxy on the left can be equally well-fitted by the ”DEV” and ”SER” profiles, with ”SER” being slightly better. For the AGN, in the middle panel, none of the models is a good profile representation, with ”EXP” being selected. This bad fitting would reflect in the quality of the photo-zs computed with algorithms that use model_flux photometry only. For each source, the flux and its residual, after removing the best model, is measured within the apertures shown in the third row. Within each aperture, all colors are computed, and the last row shows how the profiles of the colors change with the distance from the center, depending on the source type. This dependence on the size is implicitly used to compute the photo-zs correctly. For the interested reader, the importance of these features and their correlation with the redshifts are further discussed in the Appendix A.
@@
dchisq141414more details are available at https://www.legacysurvey.org/dr10/catalogs/: Difference in $\chi^{2}$ between successively more-complex model fits: PSF, REX, DEV, EXP, SER. The difference is versus no source.
@@
Figure 2: Top: Optical RGB image of an inactive galaxy (Left) and a low-redshift X-ray detected AGN (Middle) and a QSO (Right). Second row: azimuth profile of the sources, with over-plotted available model profiles (PSF, REX, EXP, DEV, and SER). Third row: the apertures within which ap_flux[band] is measured. Fourth row: the residual image after subtracting the best model. Last row: within each aperture, all colors are computed; the figures show how the profiles of the colors change with the distance from the center, depending on the source type (see more details in the main text.
@@ 3.1 Construction of the sample for training
The counterparts to the training sample are then matched in Right Ascension and Declination (max separation of 1 arcsecond) with a compilation of spectroscopic redshifts available within the eROSITA collaboration. At this stage, the compilation is a mere list of catalogs from the literature collated. It encompasses approximately 13 million celestial sources, thus including duplicates and contradicting redshift measurements (see Kluge et al., 2024, for more details). Additionally, we extend the compilation by incorporating QSOs with reliable redshifts from the low-resolution spectra in Gaia DR3 (flagQSOC =0 Gaia Collaboration et al., 2023). For some sources, multiple spectroscopic matches are available, with discordant results. This can occur, for example, due to wrong line identification in low signal-to-noise spectra. Here, we prioritize redshifts from eFEDS (Aydar et al., in prep), given that most have all been visually inspected. Alternatively, a single entry is randomly selected, provided that the difference in spectroscopic redshift between the smallest and largest redshift entry $\Delta z$ does not exceed $\Delta z$ 0.05. Objects failing to meet this criterion ( $\leq 0.3\%$ of cases) are excluded from consideration to avoid the accidental assignment of incorrect redshift values to objects in the training sample. Objects with spectroscopic matches falling within the redshift range of $z\leq$ 0.002 are assumed to be galactic and, therefore, excluded from any further consideration. The compilation is enriched by including 152 proprietary entries from SDSS-V (Kollmeier et al., in prep) (SPECPRIMARY == 1, ZWARNING == 0). Within SDSS-V, the Black Hole Mapper (BHM, Anderson et al., in prep) is one of the primary surveys, specializing in long-term, time-domain investigations of AGN and the optical characterization of eROSITA X-rays sources detected during the initial 1.5 years of the all-sky survey. This part of the training sample is represented in the histogram of Figure 3 with a solid violet line.